OpenStudy (anonymous):

f(x) = 3x - 9; g(x) = x-4 , what is the value of ( g * f ) (5)?

OpenStudy (anonymous):

g(5)*f(5)

OpenStudy (anonymous):

6*1 =6

OpenStudy (ash2326):

\[f(x)=3x-9\] \[g(x)=x-4 \] \[f(x)*g(x)=(3x-9)(x-4)\] \[(f*g)(5)=(15-9)(5-4)\] \[(f*g)(5)=6*1\]

OpenStudy (lgbasallote):

first...find f(x) x g(x) (3x - 9)(x-4) 3x^2 -21x + 36 substitute by 5 75 - 105 + 36 =6